A four-digit number is formed using the digits 1, 2, 3, 5, 8, and 9. What is the probability that the number formed is even?


1
Expert's answer
2022-02-18T05:53:00-0500

A number can be even if the last digit is 2, 8

so the probability that the number is even = 2/6 = 1/3